IF a man can covers 12 metres in one second,how many kilometres can he cover in 3 hours 45 minutes by travelling the same rate?
A)168
B)162

Answer by LinnW(1048)

First convert 3 hours 45 minutes to minutes.
3 * 60 + 45 = 180 + 45 = 225 minutes
Now convert 225 to seconds
225 * 60 = 13500 seconds
The number of seconds * 12 metres per second gives us metres travelled
13500 * 12 = 162000 metres
162000 = 162 kilometres